How many liters of a 20% alcohol solution should Maria
mix with 50 liters of a 60% alcohol solution to obtain a
30% solution?

Answer by stanbon(75887)

How many liters of a 20% alcohol solution should Maria
mix with 50 liters of a 60% alcohol solution to obtain
a 30% solution?
---------------------------
Equation:
alcohol + alcohol = alcohol
0.20x + 0.60*50 = 0.30(x+50)
Multiply thru by 100 to get:
20x + 60*50 = 30x + 30*50
10x = 30*50
x = 150 liters (amt. of 20% solution needed in the mixture)
---
Cheers,
Stan H.
=============